### v1

# NFIG: Autoregressive Image Generation with Next-Frequency Prediction

Abstract
Autoregressive models have achieved promising results in natural language processing. However, for image generation tasks, they encounter substantial challenges in effectively capturing long-range dependencies, managing computational costs, and most crucially, defining meaningful autoregressive sequences that reflect natural image hierarchies. To address these issues, we present Next-Frequency Image Generation (NFIG), a novel framework that decomposes the image generation process into multiple frequency-guided stages. Our approach first generates low-frequency components to establish global structure with fewer tokens, then progressively adds higher-frequency details, following the natural spectral hierarchy of images. This principled autoregressive sequence not only improves the quality of generated images by better capturing true causal relationships between image components, but also significantly reduces computational overhead during inference. Extensive experiments demonstrate that NFIG achieves state-of-the-art performance with fewer steps, offering a more efficient solution for image generation, with 1.25 speedup compared to VAR-d20 while achieving better performance (FID: 2.81) on the ImageNet-256 benchmark. We hope that our insight of incorporating frequency-domain knowledge to guide autoregressive sequence design will shed light on future research. We will make our code publicly available upon acceptance of the paper.
1 Introduction
The synthesis of realistic and diverse images has emerged as a fundamental challenge in computer vision. Rapid progress in this field has been propelled by the success of deep generative models, such as generative adversarial networks [6, 12, 32, 13], diffusion models [8, 41, 7, 39], and autoregressive models (AR) [35, 10, 21, 38].
Despite remarkable advances in existing methods, AR models for image generation still face several fundamental challenges.
On the one hand, due to their inherently local and sequential nature, most current AR models struggle to effectively capture long-range dependencies and global structure [3]. For example, PixelCNN [35] generates an image by predicting each pixel in a raster scanning sequence, which neglects the global image structure and relationships with distant elements.
On the other hand, the generation process is computationally intensive and time-consuming, as AR models always generate pixels or patches sequentially in a predetermined order, with each new element requiring the computation of conditional probabilities based on all previously generated content [22, 33]. For instance, ViTVQ [40] requires more than 6 seconds to generate a image over 1024 steps, making it impractical for real-time applications.
Most importantly, AR models face a fundamental challenge in defining a meaningful autoregressive sequence. Traditional AR models using raster scanning or predefined arbitrary orders fail to reflect the natural hierarchical structure and dependencies in images[34]. This improper sequence design makes it difficult for models to capture the true causal relationships between different image components, ultimately affecting the coherence and visual quality of generated outputs.
To address these limitations, recent works have explored incorporating various improvements into the generation process.
For example, Taming Transformer [5] partially addresses long-range dependency challenges through its discrete latent space and transformer architecture, but still suffers from computational inefficiency. Fast PixelCNN++ [27] speeds up generation in convolutional autoregressive models by caching hidden states to avoid redundant computation, achieving up to 183× speedups, yet it doesn’t fundamentally change the autoregressive sequence design. VAR [34] leverages the Laplacian Pyramid as a prior to guide autoregressive image generation across different resolutions, achieving improved generation quality with reduced computational overhead. However, these methods do not fully exploit the potential of natural priors inherent in raw images to guide the generation process and improve the efficiency of AR models.
In fact, the natural structure of images follows a hierarchical frequency distribution—low frequencies encode global structures while high frequencies contain local details. This organization suggests an efficient autoregressive generation sequence from low to high frequencies, aligning with visual information’s natural structure. Since low-frequency components require fewer tokens to represent, this approach enhances computational efficiency. Similar frequency-progressive principles have proven effective in diffusion models, which build from low-frequency foundations before adding higher-frequency details [26].
Building on this insight, we propose a Next-Frequency Image Generation (NFIG) framework for AR models that: (1) first generates a low-frequency image with few tokens to capture global structure; (2) then progressively adds higher-frequency components conditioned on the low-frequency foundation. This process has been shown in Figure 1. Grounded in information theory, this approach efficiently represents information across the frequency spectrum using our Frequency-guided Residual-quantized VAE (FR-VAE).
Our main contributions can be summarized as three points:
We introduce a Next-Frequency Image Generation (NFIG) framework that incorporates frequency analysis into AR image generation. To the best of our knowledge, this is the first work that guides the autoregressive process by leveraging the frequency spectrum of the image;
To demonstrate the feasibility of the NFIG paradigm, we design a Frequency-guided Residual-quantized VAE as our image tokenizer. FR-VAE separates low and high-frequency components in the representation learning process, with low frequencies encoding global structure and high frequencies preserving local details. Experiments show FR-VAE achieves a reconstruction FID of 0.85, validating its image content preservation capability;
Extensive experiments show that our proposed method achieves state-of-the-art image generation quality, reaching an FID of 2.81 with a relatively small model size. This improvement paves the way for more effective and efficient AR image generation models, making them more practical for real-world applications.
2 Related Work
2.1 Image Tokenizer
Image tokenizers, which transform continuous image data into discrete representations, have become a critical component in modern image generation systems. VQ-VAE [36] introduces vector quantization into VAE, reducing the pressure on the downstream generative model by transforming the continuous latent space into a discrete one. VQ-VAE-2 [28] extends this idea with a hierarchical framework and multi-scale codebooks, where top-level codes capture global structure, and bottom-level codes model local details, enabling higher-quality reconstruction and generation at increased resolutions.
However, VQ-VAE-based methods often suffer from codebook collapse, where only a few codebook entries are effectively used. FSQ [23] attempts to address this issue by utilizing finite scalar quantization to learn the codebook, but it does not learn a meaningful feature of images.
RQ-VAE [14] employs residual quantization with a shared codebook to enhance reconstructed image quality.
XQGAN [19] introduces feature product decomposition and a residual quantizer to enhance VQ-VAE’s performance, leading to improved image generation results.
While these approaches have made significant progress in image tokenization, they often neglect the inherent multi-scale structure of natural images, which is crucial for efficient and effective representation learning.
2.2 Autoregressive Image Generation
Autoregressive image generation has demonstrated remarkable capabilities in producing high-quality images by modeling the joint distribution of image tokens as a product of conditionals[20]. PixelCNN [35] generates images pixel by pixel in a sequential order (typically from top-left to bottom-right) using masked convolutions to ensure each pixel is conditioned only on previously generated pixels. Taming Transformer [5] introduces an autoregressive approach that generates high-resolution images by predicting the next latent patch token in a discrete compressed space learned through vector quantization. Emu3 [37] patchifies an image into a series of tokens and generates images by predicting tokens in a raster-scan manner. VAR [34] incorporates the prior knowledge of Laplacian Pyramid into transformer architecture and generates images in a next-resolution manner. MAR [17] improves the quality of generated images by replacing discrete tokens with continuous features, recognizing that autoregressive models primarily need per-token probability distributions. Infinity [17] combines autoregressive modeling with Bit-wise Modeling to enhance visual details in high-resolution image synthesis. ImageFolder [18] utilizes folded image tokens to generate high-quality images in a next-scale prediction manner, achieving superior performance.
These approaches collectively demonstrate the evolution of autoregressive image generation techniques, progressing from pixel-level prediction to more sophisticated methods involving latent spaces, hierarchical structures, and physical priors. Despite their differences in implementation, all these methods share the fundamental autoregressive principle of sequentially generating image elements conditioned on previously generated content.
3 Method
Figure 2 illustrates the main workflow for NFIG, which generates an image from low-frequency to high-frequency components, iteratively improving image quality.
3.1 Frequency-guided Residual-quantized VAE
To generate images in a frequency-aware manner, we propose a Frequency-guided Residual-quantized VAE (FR-VAE) with VQ-GAN framework. The key idea is to represent lower-frequency signals with fewer tokens and higher-frequency components with more tokens. The workflow of FR-VAE has been shown in Figure 3.
3.1.1 Frequency-guided Reconstruction
As shown in Figure 4, raw images can be decomposed into components across frequency bands: low frequencies capture global structure, high frequencies preserve details. Using Frequency-guided Decomposer and Composer, these components can be losslessly recombined for complete visual representation.
Frequency-guided Decomposer.
Given a image and a encoder , there is image latent feature and . FR-VAE decomposes into several component with different frequency by Frequency-guided Decomposer via Fast Fourier Transform (FFT) as follows:
where is the element-wise product, denotes the FFT operation, is the inverse FFT, and is the -th frequency mask used to select the desired frequency range, is the total number of frequency masks, is the component corresponding to .
Frequency-guided Composer.
Frequency-guided Composer reconstructs the raw image by interpolating different frequency components to a uniform size and merging them into a single image, as illustrated in Figure 4.
where is the interpolation function, which enables the Frequency-guided Composer to process images of different frequencies at varying resolutions.
3.1.2 Frequency-guided Residual-quantization
To efficiently represent images with minimal tokens, we implement a frequency-guided residual quantization approach that addresses information loss during downsampling. Our method progressively captures different frequency components of an image through a residual learning scheme.
Residual Token Extraction.
Given a sequence of feature maps with different dimensions , where and if , and and , we supervise the learning process using accumulated signals from the lowest frequency to the current frequency band.
The residual and representation of the -th frequency component can be computed as follows:
where is the interpolation function that upsamples
to the original feature map size, and represents the difference between the accumulated frequency components up to the -th level and the learnable features.
Vector Quantization.
In general, autoregressive models utilize the discrete tokens to generate a image. To achieve this goal, we take a simple vector quantization to transform the continuous token into discrete tokens.
We define a quantizer with a learnable codebook containing code vectors.
Using this codebook, the quantizer transforms a continuous feature map into a set of discrete tokens , where each token has the corresponds to a vector . The process of finding the optimal code representation involves:
where lookup is a function that take the index of from codebook . Due to the difficulty in learning an effective codebook, we have introduced the DINO model as an adversarial discriminator to improve codebook .
3.2 Autoregressive Image Generation
To generate images progressively from low to high frequency components, we implement a decoder-only transformer framework with Adaptive Layer Normalization (AdaLN) and block-wise causal attention [34].
3.2.1 Next-Frequency Image Prediction
Unlike conventional autoregressive image generation models that employ a ”token-by-token prediction” strategy, which often neglects spatial relationships and inherent image structure. NFIG adopts a “Coarse-to-Fine Generation” approach, first synthesizes the low-frequency components of an image, then iteratively incorporates higher-frequency details, progressively refining the generated output at each step.
The generation process for next-frequency prediction can be formulated as:
where represents the matrix of code indices utilized to encode the -th frequency component, and denotes the set of possible indices.
3.2.2 Frequency Band Division Strategy
We treat the lower frequency components as the foundation or “basis” for generating , represented by . According to information theory principles, lower-frequency signals contain less information and require fewer tokens, while higher-frequency components carry more detailed information and need more tokens for accurate representation.
Consequently, we establish an increasing scale sequence for components with increasing frequency bands . Here, denotes the maximum frequency of the entire image feature map , with . We divide the frequency bands based on their corresponding resolution as:
This frequency-guided progressive approach allows our model to capture and prioritize salient components at each stage. The method improves both computational efficiency and image quality by explicitly modeling the multi-scale frequency structure inherent in natural images.
| Type | Model | rFID | gFID | IS | Pre | Rec | #Para | #Step | Time |
| GAN | BigGAN [2] | - | 6.95 | 224.5 | 0.89 | 0.38 | 112M | 1 | – |
| GAN | GigaGAN [11] | - | 3.45 | 225.5 | 0.84 | 0.61 | 569M | 1 | – |
| GAN | StyleGan-XL [32] | - | 2.30 | 265.1 | 0.78 | 0.53 | 166M | 1 | 0.75 |
| Diff. | ADM [4] | - | 10.94 | 101.0 | 0.69 | 0.63 | 554M | 250 | 420 |
| Diff. | CDM [9] | - | 4.88 | 158.7 | – | – | – | 8100 | – |
| Diff. | LDM-4-G [30] | - | 3.60 | 247.7 | – | – | 400M | 250 | – |
| Diff. | DiT-L/2 [25] | 0.9 | 5.02 | 167.2 | 0.75 | 0.57 | 458M | 250 | 77.5 |
| Diff. | DiT-XL/2 [25] | 0.9 | 2.27 | 278.2 | 0.83 | 0.57 | 675M | 250 | 112.5 |
| Diff. | L-DiT-3B [1] | 0.9 | 2.10 | 304.4 | 0.82 | 0.60 | 3.0B | 250 | 112.5 |
| Diff. | L-DiT-7B [1] | 0.9 | 2.28 | 316.2 | 0.83 | 0.58 | 7.0B | 250 | 112.5 |
| Mask. | MaskGIT [29] | 2.28 | 6.18 | 182.1 | 0.80 | 0.51 | 227M | 8 | 1.25 |
| Mask. | RCG [16] | - | 3.49 | 215.5 | – | – | 502M | 20 | 4.75 |
| AR | VQVAE-2 [29] | 2.0 | 31.11 | 45.0 | 0.36 | 0.57 | 13.5B | 5120 | – |
| AR | VQGAN [5] | 7.94 | 18.65 | 80.4 | 0.78 | 0.26 | 227M | 256 | 47.5 |
| AR | VQGAN [5] | 7.94 | 15.78 | 74.3 | – | – | 1.4B | 256 | 60 |
| AR | ViTVQ [40] | 1.28 | 4.17 | 175.1 | – | – | 1.7B | 1024 | 60 |
| AR | ViTVQ-re [40] | 1.28 | 3.04 | 227.4 | – | – | 1.7B | 1024 | 60 |
| AR | RQTran. [15] | 1.83 | 7.55 | 134.0 | – | – | 3.8B | 68 | 52.5 |
| AR | RQTran.-re [15] | 1.83 | 3.80 | 323.7 | – | – | 3.8B | 68 | 52.5 |
| AR | VAR-d16 [34] | 0.9 | 3.55 | 274.4 | 0.84 | 0.51 | 310M | 10 | 1 |
| AR | VAR-d20 [34] | 0.9 | 2.95 | 302.6 | 0.83 | 0.56 | 600M | 10 | 1.25 |
| AR | NFIG(Ours) | 0.85 | 2.81 | 332.42 | 0.77 | 0.59 | 310M | 10 | 1 |
4 Experiment
This section introduces the experimental setup, including datasets, metrics, baselines, and implementation details. We then compare NFIG with state-of-the-art approaches on image generation tasks, followed by ablation studies and motivation verification experiments to analyze component effectiveness and design choices.
4.1 Experimental Settings
Dataset. For the purpose of our experiments, we use the ILSVRC 2012 subset of ImageNet [31], which contains 1.2 million training images, 50,000 validation images, and 100,000 test images. This subset focuses on 1,000 object categories, with each category having approximately 1,200 training images, 50 validation images, and 100 test images.
Evaluation Metrics. We adopt four metrics for quantitative evaluation: Fréchet Inception Distance (FID), Inception Score (IS), Precision (Pre), and Recall (Rec).
Baselines. We compare our method with several state-of-the-art image generation approaches, including generative adversarial networks (GAN.), diffusion models (Diff.), mask diffusion (Mask.), and autoregressive model (AR.). These methods have achieved impressive results on various image generation tasks and serve as strong baselines for comparison.
Implementation Details. We implement our model using PyTorch [24] and train it on NVIDIA H100 GPUs. For the image tokenizer, FR-VAE follows the VQGAN architecture with DINO as discriminator. The image encoder leverages pretrained weights from DINOv2-base. Our frequency residual quantizer employs multiple scaling factors across different frequency bands, resulting in a total of 680 tokens. We use a codebook size of 4096 for FR-VAE. The image generator’s backbone is based on the VAR Transformer architecture with depth 16, which enables multi-scale image prediction. We optimize our model using the Adam optimizer with a learning rate of 8e-5 and a batch size of 768. The model is trained for 350 epochs on the ImageNet dataset. At the inference stage, we set CFG 4.5 and top_k 990.
4.2 Main Results
Table 1 presents a comprehensive comparison between our approach and state-of-the-art image generative models on ImageNet . The results demonstrate that NFIG achieves superior performance within the AR model family while establishing itself as a formidable competitor among diverse generative methods across different paradigms.
AR Model Comparison. NFIG achieves the best gFID (2.81) and IS (332.42) scores, significantly outperforming other AR models. Compared to VAR-16 (gFID: 3.55, IS: 274.4), our approach reduces FID by 0.74 and improves IS by more than 21%. Additionally, the proposed approach outperforms VAR-d20, a relatively larger model, while delivering 25% faster inference speed.
Cross-family Comparison. NFIG achieves competitive performance with the best models from other families. NFIG outperforms the best mask diffusion model RCG, which has a gFID score of 3.49 and IS score of 215.5. While some GANs like StyleGAN-XL have lower gFID scores (2.30) with moderate IS (265.1), and diffusion models like DiT-L/2 show excellent gFID (2.27) and strong IS (278.2), NFIG uniquely balances both metrics at high levels (gFID: 2.81, IS: 332.42). This establishes NFIG as not only the leading AR model but also a strong competitor across all model types.
Qualitative Results. The qualitative visualization of NFIG on ImageNet is shown in Figure 5. Our model demonstrates remarkable generation capabilities across diverse ImageNet categories.
4.3 Ablation Study
To investigate the effectiveness of different components in our proposed NFIG model, we conduct a thorough ablation study on the ImageNet validation set. The results are presented in Table 2, where we evaluate the performance using rFID and gFID metrics.
We start with the baseline AR model with a sequence length of 256, which achieves an rFID of 1.62 and a gFID of 18.65.
Image Tokenizer.
To further enhance the model’s performance, we progressively introduce additional components. First, incorporating frequency-guided residual quantization into the VAR framework reduces the rFID to 1.40. Next, we integrate the DINO discriminator from VAR’s tokenizer, which substantially improves the rFID from 1.40 to 0.85.
Transformer.
We then utilize the image tokenizer (FR-VAE) to train the transformer model. Without Top_k and Classifier Free Guidance (CFG), NFIG achieves a gFID of 9.7.
The addition of Top_k sampling strategy further reduces the gFID to 6.83. Finally, incorporating CFG yields the best overall performance, maintaining the rFID at 0.85 while dramatically improving the gFID to 2.81.
Our experimental results demonstrate that the combination of FR-VAE with CFG provides optimal generation quality.
Moreover, we observe that both DINO discriminator and FR-VAE contributes significantly to improving rFID for the image tokenizer. Additionally, Top_k sampling and CFG prove essential for reducing gFID. These findings highlight the importance of discriminator guidance and conditional generation strategies in enhancing image generation quality.
| type | Method | Length | Evaluation Metric | |
| rFID | gFID | |||
| 1 | AR | 256 | 1.62 | 18.65 |
| Image Tokenizer | ||||
| 2 | + FR-Quantizer | 680 | 1.40 | – |
| 3 | + DINO-Disc | 680 | 0.85 | – |
| Generation Transformer | ||||
| 4 | + AdaLN | 680 | 0.85 | 9.7 |
| 5 | + Top_k | 680 | 0.85 | 6.83 |
| 6 | + CFG | 680 | 0.85 | 2.81 |
4.4 Motivation Verification
Effectiveness of Frequency Guidance. Similar to NFIG, VAR follows a ”coarse-to-fine” approach but differs significantly in loss computation across resolutions. VAR computes loss between different resolutions and the raw image, causing disproportionately large loss values at lower resolutions. In contrast, NFIG utilizes frequency components to guide feature learning, providing more balanced loss values throughout generation. As Figure 7 shows, this leads to dramatic variations in vector quantization loss—VAR exhibits substantially higher values across all scale factors, while NFIG maintains considerably lower loss values across all resolutions.
Frequency Distribution Analysis. The experimental results in Figure 6 demonstrate the progressive refinement of generated images and the effective capture and synthesis of multi-scale visual features by the NFIG model. The frequency spectrum visualizations reveal the model’s ability to hierarchically incorporate information from low to high frequencies, resulting in generated images with rich details and natural appearance.
5 Conclusion
This paper introduces Next-Frequency Image Generation, a novel autoregressive framework that decomposes image generation into frequency-guided stages. Our key insight leverages the hierarchical spectral distribution of natural images: low-frequency components encode global structures and long-range dependencies, while high-frequency components contain local details requiring greater information entropy. By progressively generating from low to high frequencies, the proposed method significantly outperforms existing models with comparable parameter counts, demonstrating superior quality metrics while maintaining computational efficiency. Experiments confirm that our frequency-guided approach represents an important advancement in autoregressive image synthesis.

### v2

@@
-NFIG: Autoregressive Image Generation with Next-Frequency Prediction
+NFIG: Multi-Scale Autoregressive Image Generation via Frequency Ordering
 Abstract
-Autoregressive models have achieved promising results in natural language processing.
-However, for image generation tasks, they encounter substantial challenges in effectively capturing long-range dependencies, managing computational costs, and most crucially, defining meaningful autoregressive sequences that reflect natural image hierarchies.
-To address these issues, we present Next-Frequency Image Generation (NFIG), a novel framework that decomposes the image generation process into multiple frequency-guided stages.
-Our approach first generates low-frequency components to establish global structure with fewer tokens, then progressively adds higher-frequency details, following the natural spectral hierarchy of images.
-This principled autoregressive sequence not only improves the quality of generated images by better capturing true causal relationships between image components, but also significantly reduces computational overhead during inference.
-Extensive experiments demonstrate that NFIG achieves state-of-the-art performance with fewer steps, offering a more efficient solution for image generation, with 1.25 speedup compared to VAR-d20 while achieving better performance (FID: 2.81) on the ImageNet-256 benchmark.
-We hope that our insight of incorporating frequency-domain knowledge to guide autoregressive sequence design will shed light on future research.
-We will make our code publicly available upon acceptance of the paper.
+Autoregressive models have achieved significant success in image generation.
+However, unlike the inherent hierarchical structure of image information in the spectral domain, standard autoregressive methods typically generate pixels sequentially in a fixed spatial order.
+To better leverage this spectral hierarchy, we introduce Next-Frequency Image Generation (NFIG).
+NFIG is a novel framework that decomposes the image generation process into multiple frequency-guided stages.
+NFIG aligns the generation process with the natural image structure.
+It does this by first generating low-frequency components, which efficiently capture global structure with significantly fewer tokens, and then progressively adding higher-frequency details.
+This frequency-aware paradigm offers substantial advantages: it not only improves the quality of generated images but crucially reduces inference cost by efficiently establishing global structure early on.
+Extensive experiments on the ImageNet-256 benchmark validate NFIG’s effectiveness, demonstrating superior performance (FID: 2.81) and a notable 1.25 speedup compared to the strong baseline VAR-d20.
 1 Introduction
-The synthesis of realistic and diverse images has emerged as a fundamental challenge in computer vision.
-Rapid progress in this field has been propelled by the success of deep generative models, such as generative adversarial networks [6, 12, 32, 13], diffusion models [8, 41, 7, 39], and autoregressive models (AR) [35, 10, 21, 38].
+The synthesis of images has emerged as a fundamental challenge in computer vision [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
+Rapid progress in this field has been propelled by deep generative models, such as autoregressive models (AR), Generative Adversarial Networks (GANs), and diffusion models (SD).
 Despite remarkable advances in existing methods, AR models for image generation still face several fundamental challenges.
@@
 Similar frequency-progressive principles have proven effective in diffusion models, which build from low-frequency foundations before adding higher-frequency details [20].
-Building on this insight, we propose a Next-Frequency Image Generation (NFIG) framework for AR models that: (1) first generates a low-frequency image with few tokens to capture global structure; (2) then progressively adds higher-frequency components conditioned on the low-frequency foundation.
+Motivated by this insight, we propose a Next-Frequency Image Generation (NFIG) framework for AR models that: (1) first generates a low-frequency image with few tokens to capture global structure; (2) then progressively adds higher-frequency components conditioned on the low-frequency foundation.
 This process has been shown in Figure 1.
 Grounded in information theory, this approach efficiently represents information across the frequency spectrum using our Frequency-guided Residual-quantized VAE (FR-VAE).
-Our main contributions can be summarized as three points:
+Key contributions of the NFIG framework include:
 We introduce a Next-Frequency Image Generation (NFIG) framework that incorporates frequency analysis into AR image generation.
-To the best of our knowledge, this is the first work that guides the autoregressive process by leveraging the frequency spectrum of the image;
+To our knowledge, this work is the first to guide autoregressive generation using the image’s frequency spectrum, associating low frequencies with lower resolutions and high frequencies with higher resolutions;
 To demonstrate the feasibility of the NFIG paradigm, we design a Frequency-guided Residual-quantized VAE as our image tokenizer.
@@
 Experiments show FR-VAE achieves a reconstruction FID of 0.85, validating its image content preservation capability;
-Extensive experiments show that our proposed method achieves state-of-the-art image generation quality, reaching an FID of 2.81 with a relatively small model size.
+Through extensive experimentation, we show that our approach achieves state-of-the-art image generation quality, evidenced by an FID of 2.81 from a relatively small model.
 This improvement paves the way for more effective and efficient AR image generation models, making them more practical for real-world applications.
+| Symbol | Meaning | Dimension |
+| Input image | ||
+| Reconstructed image | ||
+| Feature map from VAE encoder | ||
+| The -th frequency selection mask | ||
+| Set of -th frequency component feature maps | ||
+| Scaled feature map for -th frequency component | ||
+| Quantized representation for -th frequency component | ||
+| Cumulative signal residual through level | ||
+| The learnable codebook of FR-VAE | ||
+| The -th frequency band |
 2 Related Work
-2.1 Image Tokenizer
+Autoregressive Image Generation
+Autoregressive image generation has demonstrated remarkable capabilities in producing high-quality images by modeling the joint distribution of image tokens as a product of conditionals[21].
+PixelCNN [9] generates images sequentially, processing pixels one by one (typically top-left to bottom-right).
+It employs masked convolutions so that the generation of each pixel depends solely on the pixels already generated.
+Taming Transformer [18] introduces an autoregressive approach that generates high-resolution images by predicting the next latent patch token in a discrete compressed space learned through vector quantization.
+Emu3 [22] patchifies an image into a series of tokens and generates images by predicting tokens in a raster-scan manner.
+VAR [17] incorporates the prior knowledge of Laplacian Pyramid into transformer architecture and generates images in a next-resolution manner.
+MAR [23] improves the quality of generated images by replacing discrete tokens with continuous features, recognizing that autoregressive models primarily need per-token probability distributions.
+FAR [24] attempts to enhance MAR performance through frequency-based approaches, yet lacks critical insights into the distinctive information characteristics across different frequency bands.
+Infinity [23] combines autoregressive modeling with Bit-wise Modeling to enhance visual details in high-resolution image synthesis.
+ImageFolder [25] utilizes folded image tokens to generate high-quality images in a next-scale prediction manner, achieving superior performance.
+These approaches collectively demonstrate the evolution of autoregressive image generation techniques, progressing from pixel-level prediction to more sophisticated methods involving latent spaces, hierarchical structures, and physical priors.
+Despite their differences in implementation, all these methods share the fundamental autoregressive principle of sequentially generating image elements conditioned on previously generated content.
+Image Tokenizer
 Image tokenizers, which transform continuous image data into discrete representations, have become a critical component in modern image generation systems.
@@
 While these approaches have made significant progress in image tokenization, they often neglect the inherent multi-scale structure of natural images, which is crucial for efficient and effective representation learning.
-2.2 Autoregressive Image Generation
-Autoregressive image generation has demonstrated remarkable capabilities in producing high-quality images by modeling the joint distribution of image tokens as a product of conditionals[20].
-PixelCNN [35] generates images pixel by pixel in a sequential order (typically from top-left to bottom-right) using masked convolutions to ensure each pixel is conditioned only on previously generated pixels.
-Taming Transformer [5] introduces an autoregressive approach that generates high-resolution images by predicting the next latent patch token in a discrete compressed space learned through vector quantization.
-Emu3 [37] patchifies an image into a series of tokens and generates images by predicting tokens in a raster-scan manner.
-VAR [34] incorporates the prior knowledge of Laplacian Pyramid into transformer architecture and generates images in a next-resolution manner.
-MAR [17] improves the quality of generated images by replacing discrete tokens with continuous features, recognizing that autoregressive models primarily need per-token probability distributions.
-Infinity [17] combines autoregressive modeling with Bit-wise Modeling to enhance visual details in high-resolution image synthesis.
-ImageFolder [18] utilizes folded image tokens to generate high-quality images in a next-scale prediction manner, achieving superior performance.
-These approaches collectively demonstrate the evolution of autoregressive image generation techniques, progressing from pixel-level prediction to more sophisticated methods involving latent spaces, hierarchical structures, and physical priors.
-Despite their differences in implementation, all these methods share the fundamental autoregressive principle of sequentially generating image elements conditioned on previously generated content.
-3 Method
-Figure 2 illustrates the main workflow for NFIG, which generates an image from low-frequency to high-frequency components, iteratively improving image quality.
+3 Next-Frequency Image Generation
+To provide a comprehensive understanding of our NFIG methodology, this section delves into its intricate architectural structure.
+The essential operational sequence, illustrating the flow and interaction of the system’s key components, is clearly visualized in Figure 2.
 3.1 Frequency-guided Residual-quantized VAE
+The workflow of Frequency-guided Residual-quantized VAE has been shown in Figure 2 (a).
 To generate images in a frequency-aware manner, we propose a Frequency-guided Residual-quantized VAE (FR-VAE) with VQ-GAN framework.
 The key idea is to represent lower-frequency signals with fewer tokens and higher-frequency components with more tokens.
-The workflow of FR-VAE has been shown in Figure 3.
 3.1.1 Frequency-guided Reconstruction
-As shown in Figure 4, raw images can be decomposed into components across frequency bands: low frequencies capture global structure, high frequencies preserve details.
-Using Frequency-guided Decomposer and Composer, these components can be losslessly recombined for complete visual representation.
+As illustrated in Figure 2 (b), raw images can be decomposed into components across different frequency bands: low frequencies encode the global structure, while high frequencies retain fine details.
+Utilizing the Frequency-guided Decomposer and Composer, these components can be recombined without loss, ensuring a complete and accurate visual representation.
 Frequency-guided Decomposer.
-Given a image and a encoder , there is image latent feature and .
-FR-VAE decomposes into several component with different frequency by Frequency-guided Decomposer via Fast Fourier Transform (FFT) as follows:
-where is the element-wise product, denotes the FFT operation, is the inverse FFT, and is the -th frequency mask used to select the desired frequency range, is the total number of frequency masks, is the component corresponding to .
+Given a image and a encoder , there is image latent feature and . FR-VAE decomposes into several component with different frequency by Frequency-guided Decomposer via Fast Fourier Transform (FFT):
+Here, represents the element-wise product, signifies the FFT operation, indicates the inverse FFT, and is the -th frequency mask used to select the desired frequency range, is the total number of frequency masks, is the component corresponding to .
 Frequency-guided Composer.
-Frequency-guided Composer reconstructs the raw image by interpolating different frequency components to a uniform size and merging them into a single image, as illustrated in Figure 4.
+Frequency-guided Composer reconstructs the raw image by interpolating different frequency components to a uniform size and merging them into a single image, as illustrated in Figure 2 (b).
 where is the interpolation function, which enables the Frequency-guided Composer to process images of different frequencies at varying resolutions.
@@
 The residual and representation of the -th frequency component can be computed as follows:
-where is the interpolation function that upsamples
-to the original feature map size, and represents the difference between the accumulated frequency components up to the -th level and the learnable features.
+where is the interpolation function that upsamples to the original feature map size, and represents the difference between the accumulated frequency components up to the -th level and the learnable features.
 Vector Quantization.
@@
 The process of finding the optimal code representation involves:
-where lookup is a function that take the index of from codebook .
-Due to the difficulty in learning an effective codebook, we have introduced the DINO model as an adversarial discriminator to improve codebook .
+From , the quantized feature map and a set of discrete tokens are obtained through quantization using the codebook . Here, is a function that finds the index of the closest entry to in codebook .
 3.2 Autoregressive Image Generation
-To generate images progressively from low to high frequency components, we implement a decoder-only transformer framework with Adaptive Layer Normalization (AdaLN) and block-wise causal attention [34].
-3.2.1 Next-Frequency Image Prediction
-Unlike conventional autoregressive image generation models that employ a ”token-by-token prediction” strategy, which often neglects spatial relationships and inherent image structure.
-NFIG adopts a “Coarse-to-Fine Generation” approach, first synthesizes the low-frequency components of an image, then iteratively incorporates higher-frequency details, progressively refining the generated output at each step.
-The generation process for next-frequency prediction can be formulated as:
-where represents the matrix of code indices utilized to encode the -th frequency component, and denotes the set of possible indices.
-3.2.2 Frequency Band Division Strategy
-We treat the lower frequency components as the foundation or “basis” for generating , represented by .
+To generate images progressively from low to high frequency components, we implement a decoder-only transformer framework and block-wise causal attention [17].
+Next-Frequency Image Prediction
+Unlike conventional autoregressive image generation models that employ a "token-by-token prediction" strategy, which often neglects spatial relationships and inherent image structure.
+NFIG adopts a “Coarse-to-Fine Generation" approach, first synthesizes the low-frequency components of an image, then iteratively incorporates higher-frequency details, progressively refining the generated output at each step, as shown in Figure 2 (c).
+The generation process for next-frequency prediction is given by the autoregressive factorization:
+where is the matrix of code indices for the -th frequency component, and the set represents all available index values.
+Frequency Band Division Strategy
+We treat the lower frequency components as the foundation for generating , represented by .
 According to information theory principles, lower-frequency signals contain less information and require fewer tokens, while higher-frequency components carry more detailed information and need more tokens for accurate representation.
@@
 The method improves both computational efficiency and image quality by explicitly modeling the multi-scale frequency structure inherent in natural images.
+4 Experiment
+This section details our experimental methodology, covering datasets, evaluation metrics, comparison baselines, and implementation specifics.
+We then evaluate NFIG against state-of-the-art approaches on image generation benchmarks.
+Subsequently, ablation studies and motivation verification experiments are conducted to analyze the impact of different components and validate design decisions.
+4.1 Experimental Settings
+Dataset.
+For the purpose of our experiments, we use the ILSVRC 2012 subset of ImageNet [31], which comprises a total of 1.2 million training images, 50k validation images, and 100k test images.
+This subset focuses on 1k object categories, with each category having approximately 1.2k training images, 50 validation images, and 100 test images.
+Evaluation Metrics.
+We adopt four metrics for quantitative evaluation: Fréchet Inception Distance (FID) which measures distribution similarity between generated and real images, Inception Score (IS) which assesses quality and diversity, and Precision (Pre) and Recall (Rec) which evaluate sample fidelity and diversity coverage respectively.
+Baselines.
+Our method is benchmarked against several leading image generation techniques, including generative adversarial networks (GAN), diffusion models (Diff.), mask diffusion (Mask.), and autoregressive models (AR).
+These approaches have demonstrated strong performance on various image synthesis tasks and serve as robust comparators.
 | Type | Model | rFID | gFID | IS | Pre | Rec | #Para | #Step | Time |
@@
 | AR | NFIG(Ours) | 0.85 | 2.81 | 332.42 | 0.77 | 0.59 | 310M | 10 | 1 |
-4 Experiment
-This section introduces the experimental setup, including datasets, metrics, baselines, and implementation details.
-We then compare NFIG with state-of-the-art approaches on image generation tasks, followed by ablation studies and motivation verification experiments to analyze component effectiveness and design choices.
-4.1 Experimental Settings
-Dataset.
-For the purpose of our experiments, we use the ILSVRC 2012 subset of ImageNet [31], which contains 1.2 million training images, 50,000 validation images, and 100,000 test images.
-This subset focuses on 1,000 object categories, with each category having approximately 1,200 training images, 50 validation images, and 100 test images.
-Evaluation Metrics. We adopt four metrics for quantitative evaluation: Fréchet Inception Distance (FID), Inception Score (IS), Precision (Pre), and Recall (Rec).
-Baselines.
-We compare our method with several state-of-the-art image generation approaches, including generative adversarial networks (GAN.), diffusion models (Diff.), mask diffusion (Mask.), and autoregressive model (AR.).
-These methods have achieved impressive results on various image generation tasks and serve as strong baselines for comparison.
 Implementation Details.
-We implement our model using PyTorch [24] and train it on NVIDIA H100 GPUs.
-For the image tokenizer, FR-VAE follows the VQGAN architecture with DINO as discriminator.
-The image encoder leverages pretrained weights from DINOv2-base.
-Our frequency residual quantizer employs multiple scaling factors across different frequency bands, resulting in a total of 680 tokens.
-We use a codebook size of 4096 for FR-VAE.
-The image generator’s backbone is based on the VAR Transformer architecture with depth 16, which enables multi-scale image prediction.
-We optimize our model using the Adam optimizer with a learning rate of 8e-5 and a batch size of 768.
-The model is trained for 350 epochs on the ImageNet dataset.
-At the inference stage, we set CFG 4.5 and top_k 990.
+Our model was implemented using the PyTorch framework [42] and trained on NVIDIA H100 graphics cards.
+To ensure the reproducibility of our experiments, our implementation is built upon open-source research code, while incorporating improvements specific to this study.
+For the image tokenizer, the FR-VAE incorporates a VQGAN architecture with a DINO discriminator.
+The image encoder is initialized with pretrained weights from DINOv2-base.
+The frequency residual quantizer employs multiple scaling factors across different frequency bands, resulting in a vocabulary size of 680 tokens.
+The FR-VAE codebook size of 4096 was utilized.
+The image generator employs a VAR Transformer backbone with a depth of 16, enabling multi-scale image prediction.
+Optimization was performed using the Adam optimizer, setting the learning rate to and the batch size to 768.
+Training of the model ran for 350 epochs on the ImageNet dataset.
+For inference, we configured CFG to 4.5 and top_k to 990.
 4.2 Main Results
-Table 1 presents a comprehensive comparison between our approach and state-of-the-art image generative models on ImageNet .
-The results demonstrate that NFIG achieves superior performance within the AR model family while establishing itself as a formidable competitor among diverse generative methods across different paradigms.
+Table 2 provides a detailed comparison of our approach against leading image generative models evaluated on ImageNet .
+The findings indicate that NFIG achieves superior performance within the AR model family while establishing itself as a formidable competitor among diverse generative methods across different paradigms.
 AR Model Comparison.
@@
 This establishes NFIG as not only the leading AR model but also a strong competitor across all model types.
-Qualitative Results. The qualitative visualization of NFIG on ImageNet is shown in Figure 5. Our model demonstrates remarkable generation capabilities across diverse ImageNet categories.
+Qualitative Results. Figure 3 qualitatively shows NFIG’s impressive ability to generate diverse ImageNet images across a wide variety of categories.
 4.3 Ablation Study
-To investigate the effectiveness of different components in our proposed NFIG model, we conduct a thorough ablation study on the ImageNet validation set.
-The results are presented in Table 2, where we evaluate the performance using rFID and gFID metrics.
+To evaluate the contribution of various components within our proposed NFIG model, we perform a comprehensive ablation analysis on the ImageNet validation set.
+Table 3 summarizes the results of this study, with performance evaluated using rFID and gFID.
 We start with the baseline AR model with a sequence length of 256, which achieves an rFID of 1.62 and a gFID of 18.65.
 Image Tokenizer.
-To further enhance the model’s performance, we progressively introduce additional components.
+We incrementally add components to progressively improve the model’s performance.
 First, incorporating frequency-guided residual quantization into the VAR framework reduces the rFID to 1.40.
@@
 Transformer.
-We then utilize the image tokenizer (FR-VAE) to train the transformer model. Without Top_k and Classifier Free Guidance (CFG), NFIG achieves a gFID of 9.7.
+We then utilize the image tokenizer (FR-VAE) to train the transformer model.
+Without Top_k and Classifier Free Guidance (CFG), NFIG achieves a gFID of 9.7.
 The addition of Top_k sampling strategy further reduces the gFID to 6.83.
@@
 Additionally, Top_k sampling and CFG prove essential for reducing gFID.
-These findings highlight the importance of discriminator guidance and conditional generation strategies in enhancing image generation quality.
-| type | Method | Length | Evaluation Metric | |
+These results underscore the significance of discriminator guidance and conditional generation strategies for improving image generation quality.
+4.4 Motivation Verification
+Frequency Distribution Analysis.
+The experimental results in Figure 4 demonstrate the progressive refinement of generated images and the effective capture and synthesis of multi-scale visual features by the NFIG model.
+The frequency spectrum visualizations reveal the model’s ability to hierarchically incorporate information from low to high frequencies, resulting in generated images with rich details and natural appearance.
+These results demonstrate that NFIG’s frequency-guided approach enables more effective feature learning, particularly at lower resolutions, by maintaining balanced loss values throughout the hierarchical generation process.
+Frequency Guidance.
+Similar to NFIG, VAR follows a "coarse-to-fine" approach but differs significantly in loss computation across resolutions.
+VAR computes loss between different resolutions and the raw image, causing disproportionately large loss values at lower resolutions.
+In contrast, NFIG utilizes frequency components to guide feature learning, providing more balanced loss values throughout generation.
+As Figure 5 shows, this leads to dramatic variations in vector quantization loss—VAR exhibits substantially higher values across all scale factors, while NFIG maintains considerably lower loss values across all resolutions.
+| Type | Method | Length | Metric | |
 | rFID | gFID | |||
@@
 | 6 | + CFG | 680 | 0.85 | 2.81 |
-4.4 Motivation Verification
-Effectiveness of Frequency Guidance.
-Similar to NFIG, VAR follows a ”coarse-to-fine” approach but differs significantly in loss computation across resolutions.
-VAR computes loss between different resolutions and the raw image, causing disproportionately large loss values at lower resolutions.
-In contrast, NFIG utilizes frequency components to guide feature learning, providing more balanced loss values throughout generation.
-As Figure 7 shows, this leads to dramatic variations in vector quantization loss—VAR exhibits substantially higher values across all scale factors, while NFIG maintains considerably lower loss values across all resolutions.
-Frequency Distribution Analysis.
-The experimental results in Figure 6 demonstrate the progressive refinement of generated images and the effective capture and synthesis of multi-scale visual features by the NFIG model.
-The frequency spectrum visualizations reveal the model’s ability to hierarchically incorporate information from low to high frequencies, resulting in generated images with rich details and natural appearance.
 5 Conclusion
@@
 Experiments confirm that our frequency-guided approach represents an important advancement in autoregressive image synthesis.
+6 Limitation and Future Work
+Our frequency-guided autoregressive image generation approach shows promise, but has limitations.
+A primary issue is the simplistic frequency band division by scale, which inadequately captures information in the first band.
+Implementing a more rigorous division based on statistical analysis and physical principles would likely enhance NFIG’s performance.
+Beyond 2D spatial frequency, future work could extend to video generation, a more comprehensive direction that requires incorporating both temporal and spatial frequencies.
+Similarly, 3D object generation offers a better approximation of real-world light fields than 2D projections, where frequency analysis is vital for accurate representation.
+Due to computational constraints and time limitations, these promising directions remain unexplored and are left for future investigation.
+Appendix A: Fourier Analysis in Natural Images
+This appendix provides essential mathematical formulations and conceptual insights into the Fourier analysis of natural images, building upon the concepts discussed in the main text.
+We focus on the Discrete Fourier Transform (DFT) and its implications for image representation and the characteristics of natural scenes.
+A.1 2D Discrete Fourier Transform (2D DFT)
+The 2D DFT transforms an digital image from the spatial domain (where and are spatial coordinates) to the frequency domain, yielding an representation (where and are frequency coordinates).
+The formula is given by:
+Here, is the imaginary unit (), and the exponential term represents the basis functions (complex sinusoids) at different frequencies .
+A.2 Inverse 2D Discrete Fourier Transform (2D IDFT)
+The 2D IDFT allows us to reconstruct the original spatial domain image from its frequency domain representation . The formula is:
+Note the scaling factor and the positive sign in the exponent compared to the forward transform.
+A.3 Magnitude Spectrum and Power Spectrum
+The frequency domain representation obtained from the DFT is generally a complex number.
+Its magnitude, , is known as the Magnitude Spectrum, which quantifies the amplitude of each frequency component present in the image.
+Closely related is the Power Spectrum (or Power Spectral Density), defined as the square of the magnitude spectrum.
+It represents how the total signal energy is distributed across the different frequencies:
+A key characteristic of natural images is that their power spectrum typically exhibits a rapid decay as frequency increases.
+Specifically, the power tends to fall off with increasing radial frequency , often approximated by a law, where is a constant typically around 2.
+This 1/f property implies that low spatial frequencies (corresponding to coarse structures and overall variations) contain significantly more energy than high spatial frequencies (corresponding to fine details and sharp transitions).
+This fundamental statistical feature of natural images is widely utilized and modeled in various image processing and computer vision tasks.
+Appendix B: Conceptual Parallel: Quantum Field Theory and Frequency-Based Image Generation
+A Conceptual Parallel: Quantum Field Theory and Frequency-Based Image Generation – An Inspiring Perspective
+Introduction
+Quantum Field Theory (QFT) provides a fundamental description of the universe, viewing physical reality in terms of quantum fields rather than discrete point particles.
+Interestingly, this framework, particularly its treatment of fields and their excitations, presents some insightful conceptual parallels with modern frequency-based approaches to image generation, such as NFIG.
+Exploring this connection might offer new perspectives for both fields.
+Core Concepts and Parallels
+At its core, QFT sees the universe as composed of continuous quantum fields that fill all of spacetime.
+Particles, like electrons or photons, emerge as quantized excitations of these underlying fields.
+A key aspect is that these fields can be described as a collection of fundamental "modes," each characterized by specific energy and momentum.
+These quantities are linked to frequency and wavevector through the fundamental Planck-Einstein and de Broglie relations:
+where is the reduced Planck constant, is angular frequency, and is the wavevector.
+Thus, exciting a specific mode of the quantum field corresponds to creating a particle with a definite frequency and spatial variation.
+Correspondingly, frequency-based image generation methods view an image as a composition of different spatial frequency components.
+Techniques like NFIG often build images hierarchically, starting with lower-frequency content to establish the coarse structure and overall composition, and then progressively adding higher-frequency components to introduce fine details.
+Drawing on this, we can identify several conceptual parallels.
+Both QFT and frequency-based image generation involve the idea of decomposing a complex system (physical reality or an image) into fundamental constituents characterized by frequency or momentum/wavevector.
+In QFT, physical states are understood as arising from exciting field modes, while in image generation, images are built from spatial frequency components.
+Furthermore, both often involve a hierarchical structure: QFT relates different energy/momentum scales to distinct physical phenomena and interactions, while frequency bands in image generation capture image features at different spatial scales, from coarse structures (low frequencies) to fine details (high frequencies).
+Finally, just as complex quantum states are formed by superposing mode excitations, a complete image is formed by superimposing its various frequency components.
+Potential Research Insights
+Nevertheless, exploring this conceptual link might stimulate thinking and offer potential avenues for future research.
+One might consider whether mathematical structures or principles from QFT could inform the design of new image generation algorithms, particularly for multi-scale or frequency-space operations.
+Similarly, QFT’s approach to relating different energy scales to physical phenomena could potentially provide inspiration for learning more effective multi-scale representations within neural networks.
+More broadly, this perspective hints at the possibility of developing more unified theoretical frameworks that could potentially bridge concepts in fundamental physics and complex learned systems like neural networks.
+Conclusion
+Ultimately, this conceptual parallel serves as an inspiring thought framework rather than a strict equivalence.
+It highlights shared principles in how complex systems can be understood and constructed through their fundamental frequency/momentum components, potentially fostering interdisciplinary dialogue and novel approaches.